# Nonequilibrium steady states in multi-bath quantum collision models

### v1

@@ III.2 Steady State Correlations
We now consider the behavior of system-environment correlations. To do this we must employ a suitable measure and, for our purposes, we choose the negativity based on the postive partial transpose criterion. For a two qubit system, it is defined Peres (1996),
@@
where $\lambda_{neg,j}^{A(BC)}$ are the negative eigenvalues of the partially transposed density matrix of the bipartition $A$-$BC$. The genuine tripartite entanglement of the three-qubit state can then be calculated by taking a geometric average of the three possible bipartitions,
@@
In Fig. 5 we examine the entanglement established in the steady state of setting II in the full tripartite state as well as the various reduced state bipartitions. In all cases we find that there is a vanishingly small degree of entanglement present for sufficiently fast system-environment collisions and weak intra-bath interactions, i.e. small $\Delta t$ and $\delta$. Increasing these parameters allows for the establishment of entanglement in the tripartite system as shown in panel (a). The region of non-zero entanglement closely aligns with regions where setting II’s steady state significantly deviates from the expected thermal state. Considering the reduced state of the system and one of the baths, shown in panels (b) and (c), we again see a good agreement between regions where the system’s steady state is at a different effective temperature and the presence of bipartite entanglement between $S$ and $A_{n}^{(i)}$. For the reduced state between the two bath units, shown in panel (d), we see that entanglement is only established for large values of both $\Delta t$ and $\delta$. Interestingly, this region corresponds to the regions where $\Delta\beta$ is positive (see Fig. 4(d)), i.e. the establishment of cross-environmental entanglement aligns with the system being driven to cooler effective temperature.
@@
Figure 5: Steady state entanglement arising from setting II. Panel (a): Genuine tripartite entanglement between the system and two “memory” units. Panels (b-d): Bipartite entanglement of the various reduced states. In all panels we fix $\beta\!=\!2,\omega=1,\Gamma=1$.
It is worth reminding that the above is in stark contrast to setting I’s steady state, which is driven to a factorised state between S and $A_{n}$ and therefore clearly there is no entanglement present in the steady state Cusumano et al. (2018). This can be intuitively understood due to the fact that for the considered interactions, setting I leads to homogenization and therefore in the steady state the system and all bath auxiliary units are all in identical states and, as the collisions correspond to a partial swap, they do not change the states. As demonstrated by Fig. 5, this is no longer the case in setting II, where the system evidently cannot homogenize and therefore each collisions is, generally, a strongly entangling operation.
We note that the entanglement in the tripartite NESS of system and memories is distinct from entanglement that is created during each collision of the memories with the incoming bath auxiliaries. Indeed, even in the Markovian case, once the system has reached the steady state, each subsequent collision in setting II will lead to the creation of genuine tripartite entanglement between the system and the bath units while leaving the reduced state of the system unchanged. This implies that the stronger deviation form the expected thermal state in the non-Markovian case is a consequence of setting II’s ability to create entanglement in the system memory compound in addition to the entanglement being continually created and destroyed in each subsequent collision with the baths. We therefore conclude that the presence of non-Markovianity enhances the non-equilibrium character of the state due to the map’s ability to create entanglement. However, we remark that as setting I has a comparable degree of non-Markovianity present in the dynamics (see Fig. 3), it follows that its presence is not solely responsible for the observed deviation from the expected thermal state and rather it is a more complex interplay between the constant energy exchanges, the creation and maintenance of strong quantum correlations, and the non-Markovian character of the dynamics.
@@ IV Stochastic Heat Currents from Trajectories
Quantum collision models lend themselves naturally to the simulation of quantum trajectories when we consider performing measurements on the auxiliary units immediately after their interaction with the system Brun (2002). It is known that even for non-Markovian collision models, individual trajectories can be averaged to recover the unconditional dynamics Kretschmer et al. (2016b); Whalen (2019). In this section we will show how the unconditional heat current can be recovered from two-point measurements on the auxiliary qubits. The stochastic heat current after one complete collision is calculated by performing a projective measurement, with measurement operator $P_{j}$, in the energy eigen-basis of the auxiliary unit. This measurement is performed twice, first before the auxilliary unit has undergone any interaction with any other system, and again after it has completed all of its interactions. This procedure is commonly referred to as a two-point measurement (TPM) scheme. The post-measurement states of the auxiliary unit after the first and second measurements are given by,
@@
where $\tilde{\rho}^{\prime}_{A_{n}^{(i)}}$ is the state of the auxiliary unit after it has undergone the first measurement and been involved in the the $A^{(i)}_{n-1}$-$A^{(i)}_{n}$,z $S$-$A^{(i)}_{n}$ and $A^{(i)}_{n}$-$A^{(i)}_{n+1}$ collisions. The stochastic heat current for the $n$-th time step is then
@@
where the averaging is done over $M$ trajectories. To give an illustrative example, in Fig. 6 we calculate the stochastic heat current for the zero temperature bath in setting II for two different values of $M$ and compare it to the unconditional heat current calculated using Eq. (13). It should be noted that since in this setting the zero temperature bath starts in the |$\ket{0}$ state, the first measurement in the TPM scheme has no effect and the TPM measurement scheme is equivalent to a typical continuous monitoring in this set-up. We see close agreement between the averaged stochastic heat and the unconditional value, with this agreement improving with increasing $M$.
Finally, it is interesting to note that certain heat fluxes in the non-Markovian setting, specifically those associated with the $S$-$A_{n}$ collisions, cannot be accessed using this type of TPM scheme. For example, in order to access the heat flux specific to the $S$-$A_{n}^{(0)}$ interaction a measurement must be performed on $A_{n}^{(0)}$ after the $S$-$A_{n}$ interaction, but before the $A_{n}^{(0)}$-$A_{n+1}^{(0)}$ interaction takes place. Clearly such a measurement can be done, however it will clearly affect the correlations shared between the $S$ and $A_{n+1}^{(0)}$ as well as the incoming state of $A_{n+1}^{(0)}$. This leads to a significant change in system dynamics and thus we will have a different stochastic head current compared to the unconditional dynamics. These “inaccessible currents” appear to be a unique feature of the non-Markovian setting.
Figure 6: Mean stochastic heat current into the zero temperature bath, $\langle\Delta\tilde{\mathcal{Q}}_{n}^{(0)}\rangle_{M}$ against $t\!=\!n\Delta t$ for $M=10,000$ (green, dashed line) and $M=100,000$ (black line) along with the corresponding unconditional heat current $\Delta\mathcal{Q}^{(0)}_{n}$ (magenta line). In all cases the system is initialized in the ground state $\rho_{0}=|0\rangle\langle 0|$. Parameters used: $\delta=0.95\frac{\pi}{2},\Gamma=1,\beta=1$.
@@ V Conclusions
We have examined the steady state properties of a two-level quantum system undergoing an open dynamics simulated using quantum collision models. In the Markovian case, we considered a simple thermalisation process which can be effectively modelled via a standard GKSL master equation and considered two settings that precisely match this dynamics in the limit of vanishing collision time, $\Delta t\!\to\!0$.
@@
As a benchmark we considered the typical situation, referred to as setting I, where the collisional bath consists of identical auxiliary units all prepared in the same thermal state. For the considered system-environment interactions, setting I is known to lead to thermalisation (and more generally homogenisation), leaving the system in equilibrium with the bath and, furthermore, the system and environment share no correlations in the steady state. By contrast, setting II involves two collisional baths, whose auxiliary qubits are prepared in either their ground or excited states, and the bath temperature is encoded through the system-bath couplings. As a main result we have demonstrated that despite the dynamics of the system being formally identical in the $\Delta t\!\to\!0$ limit, setting II necessarily drives the system to a nonequilibrium steady state (NESS), supported by non-zero steady state heat currents. By exploiting the flexibility of collision models, we examined the impact of non-Markovian dynamics on the resulting steady state, by introducing additional intra-environment interactions. Despite setting I and setting II exhibiting similar amounts of non-Markovianity, as quantified via the information back flow, we showed that the non-Markovian case of setting II results in the system being driven further from the expected equilibrium state. Although setting I always reaches the canonical temperature of the bath in the limit $\Delta t\to 0$, the non-equilibrium steady state of setting II can reach a different temperature. Through a careful analysis of the tripartite and bipartite entanglement, we have established that the non-equilibrium character of the steady state is principally rooted in the ability of setting II to create and maintain strong quantum correlations. We have also demonstrated how the average heat current can be recovered in setting II using a two-point measurement scheme, even in the presence of non-Markovianity.
@@
Our results provide insight into the dynamics and thermodynamics of thermalisation and equilibration and highlight that seemingly equivalent dynamics at the level of the system of interest can correspond to fundamentally different physical processes. The versatility of the collision model framework allows them to be employed in the modeling of practical quantum devices and offers a promising avenue for future work. One possible direction is to model non-Markovian fermionic environments relevant to semiconductor quantum dot devices coupled to electronic circuits Goldhaber-Gordon et al. (1998); Shim et al. (2018); Barthelemy and Vandersypen (2013); Iftikhar et al. (2018); Pouse et al. (2023); Petropoulos et al. (2024), where we expect memory effects and strong system-bath correlations to play an important role. This builds on recent work showing that non-Markovian collision models are a powerful and systematic method for treating bosonic environments Lacroix et al. (2024). Another possibility would be to use the framework of non-Markovian collisional trajectories for measurement-based feedback control in non-Markovian environments, for example, to see the effects of specific intra-environment interactions and the resulting environmental memory on the ability to prepare a given target state Harwood et al. (2023).

### v2

@@ III.2 Steady State Correlations
We now consider the behavior of system-environment correlations. To do this we must employ a suitable measure and, for our purposes, we choose the negativity based on the postive partial transpose criterion. For a two qubit system, it is defined BiPartite_Neg,
@@
where $\lambda_{neg,j}^{A(BC)}$ are the negative eigenvalues of the partially transposed density matrix of the bipartition AA-BCBC. The genuine tripartite entanglement of the three-qubit state can then be calculated by taking a geometric average of the three possible bipartitions sabin2008classification,
@@
In Fig. 5 we examine the entanglement established in the steady state of setting II in the full tripartite state as well as the various reduced state bipartitions. In all cases we find that there is a vanishingly small degree of entanglement present for sufficiently fast system-environment collisions and weak intra-environment interactions, i.e. small $\Delta t$ and $\delta$. Increasing these parameters allows for the establishment of entanglement in the tripartite system as shown in panel (a). The region of non-zero entanglement closely aligns with regions where setting II’s steady state significantly deviates from the expected thermal state. Considering the reduced state of the system and one of the baths, shown in panels (b) and (c), we again see a good agreement between regions where the system’s steady state is at a different effective temperature and the presence of bipartite entanglement between $S$ and $A_{n}^{(i)}$. For the reduced state between the two bath units, shown in panel (d), we see that entanglement can be established, however, it becomes appreciable only for relatively large values of both $\Delta t$ and $\delta$.
@@
Figure 5: Steady state entanglement arising from setting II. Panel (a): Genuine tripartite entanglement between the system and two “memory” units. Panels (b-d): Bipartite entanglement of the various reduced states. In all panels we fix $\beta\!=\!2,\omega=1,\Gamma=4$.
It is worth reminding that the above is in stark contrast to setting I’s steady state, which is driven to a factorised state between $S$ and $A_{n}$ and therefore clearly there is no entanglement present in the steady state GiovannettiPRA2018. This can be intuitively understood due to the fact that for the considered interactions, setting I leads to homogenization and therefore in the steady state the system and all bath auxiliary units are all in identical states and, as the intra-environment collisions correspond to a partial swap, they do not change the states. As demonstrated by Fig. 5, this is no longer the case in setting II, where the system evidently cannot homogenize and therefore each collisions is, generally, a strongly entangling operation.
We note that the entanglement in the tripartite NESS of system and memories is distinct from entanglement that is created during each collision of the memories with the incoming bath auxiliaries. Indeed, even in the Markovian case, once the system has reached the steady state, each subsequent collision in setting II will lead to the creation of genuine tripartite entanglement between the system and the bath units while leaving the reduced state of the system unchanged. This implies that the deviation from the expected thermal state in the non-Markovian case is a consequence of setting II’s ability to create entanglement in the system memory compound in addition to the entanglement being continually created and destroyed in each subsequent collision with the baths. We therefore conclude that the presence of non-Markovianity enhances the non-equilibrium character of the state due to the map’s ability to create entanglement. However, we remark that as setting I has a comparable degree of non-Markovianity present in the dynamics (see Fig. 3), it follows that its presence is not solely responsible for the observed deviation from the expected thermal state and rather it is a more complex interplay between the constant energy exchanges, the creation and maintenance of strong quantum correlations, and the non-Markovian character of the dynamics.
@@ IV Stochastic Heat Currents from Trajectories
Quantum collision models lend themselves naturally to the simulation of quantum trajectories when we consider performing measurements on the auxiliary units immediately after their interaction with the system Brun_CM_Trajectories. It is known that even for non-Markovian collision models, individual trajectories can be averaged to recover the unconditional dynamics Kretschmer_CompositeCM; Whalen_NMCM_Trajectories; cilluffo2021microscopic. In this section we will show how the unconditional heat current can be recovered from two-point measurements on the auxiliary qubits. The stochastic heat current after one complete collision is calculated by performing a projective measurement, with measurement operator $P_{j}$, in the energy eigen-basis of the auxiliary unit. This measurement is performed twice, first before the auxilliary unit has undergone any interaction with any other system, and again after it has completed all of its interactions. This procedure is commonly referred to as a two-point measurement (TPM) scheme. The post-measurement states of the auxiliary unit after the first and second measurements are given by,
@@
where $\tilde{\rho}^{\prime}_{A_{n}^{(i)}}$ is the state of the auxiliary unit after it has undergone the first measurement and been involved in the the $A^{(i)}_{n-1}-$$A^{(i)}_{n}$, $S$-$A^{(i)}_{n}$ and $A^{(i)}_{n}-$$A^{(i)}_{n+1}$ collisions. The stochastic heat current for the $n$-th time step is then
@@
[…]
Finally, it is interesting to note that certain heat fluxes in the non-Markovian setting, specifically those associated with the $S$-$A_{n}$ collisions, cannot be accessed using this type of TPM scheme. For example, in order to access the heat flux specific to the $S$-$A_{n}^{(0)}$ interaction a measurement must be performed on $A_{n}^{(0)}$ after the $S$-$A_{n}$ interaction, but before the $A_{n}^{(0)}$-$A_{n+1}^{(0)}$ interaction takes place. While such a measurement can be done, clearly it will affect the correlations shared between the $S$ and $A_{n+1}^{(0)}$ as well as the incoming state of $A_{n+1}^{(0)}$. This leads to a significant change in system dynamics and thus we will have a different stochastic heat current compared to the unconditional dynamics. These “inaccessible currents” appear to be a unique feature of the non-Markovian setting.
Figure 6: Mean stochastic heat current into the zero temperature bath, $\langle\Delta\tilde{\mathcal{Q}}_{n}^{(0)}\rangle_{M}$ against $t\!=\!n\Delta t$ for $M=10,000$ (green, dashed line) and $M=100,000$ (black line) along with the corresponding unconditional heat current $\Delta\mathcal{Q}^{(0)}_{n}$ (magenta line). In all cases the system is initialized in the ground state $\rho_{0}=|0\rangle\langle 0|$. Parameters used: $\delta=0.95\frac{\pi}{2},\Gamma=4,\beta=1$, $\Delta t=0.0025$.
@@ V Conclusions
We have examined the steady state properties of a two-level quantum system undergoing an open dynamics simulated using quantum collision models. In the Markovian case, we considered a simple thermalisation process which can be effectively modelled via a standard GKSL master equation and considered two settings that drive the system to the same steady state.
@@
As a benchmark we considered the typical situation, referred to as setting I, where the collisional bath consists of identical auxiliary units all prepared in the same thermal state. For the system-environment interactions given by Eq. (6), setting I is known to lead to thermalisation, leaving the system in equilibrium with the bath and, furthermore, the system and environment share no correlations in the steady state. By contrast, setting II involves two collisional baths, whose auxiliary qubits are prepared in either their ground or excited states, and the bath temperature is encoded through the system-bath couplings. As a main result we have demonstrated that despite both settings recovering the GKSL master equation for the system in the $\Delta t\!\to\!0$ limit, setting II necessarily drives the system to a nonequilibrium steady state (NESS), supported by non-zero steady state heat currents. By exploiting the flexibility of collision models, we examined the impact of non-Markovian dynamics on the resulting steady state, by introducing additional intra-environment interactions. Although setting I and setting II exhibit similar amounts of non-Markovianity, as quantified via the information back flow, we showed that the non-Markovian case of setting II results in the system being driven away from the expected thermal state. Although setting I always reaches the canonical temperature of the bath, the non-equilibrium steady state of setting II can reach a different temperature. Through a careful analysis of the tripartite and bipartite entanglement, we have established that the non-equilibrium character of the steady state is rooted in the ability of setting II to create and maintain strong quantum correlations. We have also demonstrated how the average heat current can be recovered in setting II using a two-point measurement scheme, even in the presence of non-Markovianity.
@@
Our results provide insight into the dynamics and thermodynamics of thermalisation and equilibration and highlight that seemingly equivalent dynamics at the level of the system of interest can correspond to fundamentally different physical processes. The versatility of the collision model framework allows them to be employed in the modeling of practical quantum devices and offers a promising avenue for future work. One possible direction is to model non-Markovian fermionic environments relevant to semiconductor quantum dot devices coupled to electronic circuits goldhaber1998kondo; shim2018numerical; barthelemy2013quantum; iftikhar2018tunable; pouse2023quantum; petropoulos2024nanoscale, where we expect memory effects and strong system-bath correlations to play an important role. This builds on recent work showing that non-Markovian collision models are a powerful and systematic method for treating bosonic environments Lacroix2024MakingQCMExact. Another possibility would be to use the framework of non-Markovian collisional trajectories for measurement-based feedback control in non-Markovian environments, for example, to see the effects of specific intra-environment interactions and the resulting environmental memory on the ability to prepare a given target state HarwoodCMFeedback.
Appendix
V.1 Continuous-time limit of setting II
Following similar steps as Refs. CICCARELLO20221; cusumano2022quantum we can show that GKSL master equation, Eq. (4) is recovered in the time-continuous limit also for setting II. We assume an arbitrary form for the system-auxiliary interaction
$H_{I}=\sum_{\nu}g_{\nu}A_{\nu}B_{\nu}$ (23)
where $g_{\nu}$ are coefficients, $A_{\nu}$ and $B_{\nu}$ are sets of system and auxiliary operators, respectively, and $\Delta\rho_{n}=\rho_{n}-\rho_{n-1}$. By expanding the unitary operator governing a single collision up to second order in $\Delta t$, it can be readily shown that a finite difference equation for the state of the system can be derived CICCARELLO20221
$\frac{\Delta\rho_{n}}{\Delta t}=-i[H_{S}+\sum_{\nu}g_{\nu}\langle B_{\nu}\rangle A_{\nu},\rho_{n-1}]\\ +\sum_{\nu\mu}g_{\nu}g_{\mu}\Delta t\langle B_{\mu}B_{\nu}\rangle(A_{\nu}\rho_{n-1}A_{\mu}-\frac{1}{2}\{A_{\mu}A_{\nu},\rho_{n-1}\}$ (24)
To arrive at the master equation for the system it therefore follows that we require the first and second moments of the auxilliary operators
$\langle B_{\nu}\rangle=\text{Tr}[B_{\nu}\eta_{n}],\qquad\langle B_{\nu}B_{\mu}\rangle=\text{Tr}[B_{\nu}B_{\mu}\eta_{n}].$ (25)
where we have introduced the state of the $n$-th auxiliary, $\eta_{n}$. Note that thus far we have made no assumptions about the specific form of interaction or the structure of the auxiliary. For setting II, we assume the auxiliary is $\eta_{n}\!=\!(|0\rangle\langle 0|\otimes|1\rangle\langle 1|)$ and for convenience we rewrite the interaction as
$H_{I}=\sum_{i=0,1}-J^{(i)}(\sigma_{S}^{+}\otimes\sigma_{A^{(i)}_{n}}^{-}+\sigma_{S}^{-}\otimes\sigma_{A^{(i)}_{n}}^{+})$. (26)
It then follows that we need only compute the relevant moments with,
$A_{\nu}$ $\in\{\sigma_{S}^{+},\sigma_{S}^{-},\sigma_{S}^{+},\sigma_{S}^{-}\}$
$B_{\nu}$ $\in\{\sigma_{A_{n}^{(0)}}^{-},\sigma_{A_{n}^{(0)}}^{+},\sigma_{A_{n}^{(1)}}^{-},\sigma_{A_{n}^{(1)}}^{+}\}$
$\ g_{\nu}$ $\in\{-J^{(0)},-J^{(0)},-J^{(1)},-J^{(1)}\}.$
For this set of operators and specific form of $\eta_{n}$, we find $\sum_{\nu}g_{\nu}\langle B_{\nu}\rangle A_{\nu}=0$, while the only non-zero second moments are $\langle\sigma_{A_{n}^{(0)}}^{-}\sigma_{A_{n}^{(0)}}^{+}\rangle=\langle\sigma_{A_{n}^{(1)}}^{+}\sigma_{A_{n}^{(1)}}^{-}\rangle=1$. Equation (24) then reduces to,
$\frac{\Delta\rho_{n}}{\Delta t}=-i[H_{S},\rho_{n-1}]+(J^{(0)})^{2}\Delta t\mathcal{D}[\sigma_{S}^{-}]\rho_{n-1}\\ +(J^{(1)})^{2}\Delta t\mathcal{D}[\sigma_{S}^{+}]\rho_{n-1}.$ (27)
In order for the dissipators to survive the continuous time limit we introduce diverging coupling strengths as given in the main text
$J^{(0)}=\sqrt{\frac{\Gamma(\bar{N}+1)}{\Delta t}},\quad J^{(1)}=\sqrt{\frac{\Gamma\bar{N}}{\Delta t}},$ (28)
such that after taking $\Delta t\rightarrow 0$ we obtain,
$\dfrac{d\rho}{dt}=-i\left[H_{S},\rho\right]+\Gamma(\bar{N}+1)\mathcal{D}[\sigma^{-}]\rho+\Gamma\bar{N}\mathcal{D}[\sigma^{+}]\rho\;,$ (29)
i.e. the GKSL master equation given in Eq. (4).
V.2 Heisenberg system-environment interaction
Here we consider how changing the interaction between system and environment impacts the resulting steady states arising from the two collision model settings. In particular, we consider an isotropic Heisenberg interaction
which gives rise to precisely the partial-SWAP operation
similar to Eq. (10) in the main text. We remark that, at the level of the steady states, many of the main insights discussed in the main text are unaffected. We therefore focus on only the aspects where differences emerge.
For collision models following setting I, this type of system-environment interaction has been extensively considered ScaraniPRL2002; ruariMauro; Bassano. (One can also consider an incoherent SWAP operation, see for example Ref. karpat2025transient.). For setting I, it should be evident that this choice of interaction has no impact on the steady state that the system is driven to; the dynamics will invariably homogenize the system with initial state of the auxiliary units in both the Markovian and non-Markovian regimes. For auxiliaries again initialised in thermal states, i.e. Eq. (5), it known that despite the dynamics driving the system to the same thermal state, for this type of interaction we no longer recover the dynamics governed by Eq. (4cusumano2022quantum. Nevertheless, as our focus is on the steady-state properties our interest lies in examining the how this choice of interaction reveals more substantial differences between settings I and II.
An immediate and remarkable difference emerges even in the Markovian case; i.e. $\delta\!=\!0$. The analytical form of the steady state in setting II is readily accessible, although too cumbersome to provide any direct insight. Nevertheless, it is diagonal in the system’s energy eigenbasis and, since we are considering a simple two-level system, it can therefore be assigned an effective temperature, $\beta_{e}$, via Eq. (17). While the explicit formula of $g_{e}$ or $\beta_{e}$ is too cumbersome, to illustrate the idea, in the case of very small $\Delta t$ we have
This demonstrates that for system environment interactions governed by the partial-SWAP, even in the Markovian limit setting II drives the system to a steady state that deviates from the equilibrium value $g=(2\bar{N}+1)^{-1}$ by a term that is proportional to $\Delta t$.
(b)                        (c)
Refer to caption Refer to caption
Figure 7: (a) Steady states in the Markovian limit for system environment interactions given by Eq. (30). Main panel: Deviation of the effective temperature from the canonical temperature, $\Delta\beta$, with increasing collision time-step $\Delta t$ for different canonical temperatures $\beta\!=\!2$ (dashed, blue; lower temperature) and $\beta\!=\!0.5$ (solid, red; higher temperature). Inset: Non-equilibrium steady state heat current, $\Delta\mathcal{Q}/\Delta t$ for the initially low energy (cold) bath, $A_{n}^{(0)}$. (b) and (c) Difference in system temperature reached at the steady state, $\Delta\beta\!=\!{\beta_{e}-\beta}$ considering the impact of increasing the A-A interactions $\delta$ and increasing the S-A collision duration $\Delta t$. Other parameters used: $\omega\!=\!1,\Gamma\!=\!1$.
[…]
In panels (b) and (c) we introduce non-Markovianity via the same intra-environment interactions, Eq. (10). While it is clear that the qualitative behavior is consistent with the exchange interaction considered in the main text, we do see some differences worth highlighting. In particular we see that $\Delta\beta$ can be positive or negative depending on interplay between the strength of the intra-environment interactions and the system environment collision duration. This is markedly different from the Markovian case shown in Fig. 7(a) where $\Delta\beta\!>\!0$. This variation in the effective temperature that the steady state reaches, corresponding to a transition from effective heating to cooling, implies that in the non-Markovian two bath setting the system can still reach the expected thermal state, shown by the white dashed lines in Fig. 7(b) and (c) indicating the parameter values where $\Delta\beta\!=\!0$. This provides further evidence that while the non-Markovianity of the dynamics is a contributing factor for the system reaching a non-equilibrium steady state, there is no simple relationship between the degree of non-Markovianity and the effective temperature to which the system settles.
As a final remark, we note that in the Markovian case $\delta=0$, Eq. (32) shows that in the limit $\Delta t\to 0$ the steady state of setting II is the canonical Gibbs state with $g_{e}\to g$ and hence $\beta_{e}\to\beta$. Interestingly, in the non-Markovian case, a different effective temperature $\beta_{e}\neq\beta$ can be reached even in the limit $\Delta t\to 0$. As can be appreciated visually from Fig. 7(b) and (c), the blue regions of large temperature deviation are pushed to larger and larger $\delta$ as $\Delta t$ becomes smaller. We therefore have a subtle order-of-limits issue, which can be analyzed by considering the effective temperature of the NESS as we send both $\Delta t\to 0$ and $\delta\to 1$ while keeping their ratio $r=\Delta t/(1-\delta)$ constant and finite. By recasting the problem in terms of variables $r$ and $\Delta t$ we find numerically that substantial deviations $\Delta\beta$ survive the $\Delta t\to 0$ limit for fixed values of $r>1$. However, we were not able to extract the non-Markovian analogue of Eq. (32) analytically. Our analysis shows that non-Markovian interactions have a significant effect on the dynamics in setting II, leading to a NESS with with a temperature different from the canonical temperature, even as $\Delta t\to 0$.